The. Hypotenuse of a right triangle is 26cm. The sum of the other two sides is 34cm. Find the lengths of the other sides.
I think I am supposed to use Pythagorean but am not really sure

Answer by JulietG(1812)

In a right triangle, a^2 + b^2 = c^
Here's what we know:
c = 26
a + b = 34
.
a^2 + b^2 = 676
.
Do you have squares memorized to 25?
If so, the answer is obvious right away.
24^2 = 576. Add 10^2
.
It's a picture for me, so there were no intermediate steps (although they do exist).